OpenStudy (anonymous):

what is the derivitive of 5^x^2

OpenStudy (anonymous):

(5^(x^2))

OpenStudy (anonymous):

y = 5^(x^2) lny = x^2 * ln5 1/y * dy/dx = 2x * ln5 + x^2 * 0 dy/dx = y(2xln5) dy/dx = 5^(x^2) * 2x * ln5

OpenStudy (anonymous):

amirite?

OpenStudy (anonymous):

\[y'=5^{2x} \] \[y'=10x ^{2x-1}\]

OpenStudy (anonymous):

you have a choice. you can take the log, differentiate, and then multiply by the original function. or you can remember that \[\frac{d}{dx}b^x=b^x\ln(b)\] and use the chain rule

OpenStudy (anonymous):

ya i was thinking b^x ln(b)

OpenStudy (anonymous):

if you remember that, then this one is easy just don't forget the chain rule you get \[2x\ln(5)\times 5^{x^2}\] in one step